OpenStudy (anonymous):

How can I factor 2(x+3)^-4/3 + (x+3)^-1/3 @SolomonZelman

OpenStudy (jdoe0001):

\(\large \bf { 2(x+3)^{-\frac{4}{3}}+(x+3)^{-\frac{1}{3}} \\ \quad \\ 2(x+3)^{-\frac{1}{3}-\frac{3}{3}}+(x+3)^{-\frac{1}{3}} }\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2(x+3)^{-4/3}+(x+3)^{-1/3} }\) \(\LARGE\color{blue}{ 2\frac{1}{\sqrt[3]{(x+3)^4}}+\frac{1}{\sqrt[3]{(x+3)}} }\) \(\LARGE\color{blue}{ 2(x+3)\frac{1}{\sqrt[3]{(x+3)}}+\frac{1}{\sqrt[3]{(x+3)}} }\) \(\LARGE\color{blue}{ 2(x+3)\frac{1}{\sqrt[3]{(x+3)}}+1\frac{1}{\sqrt[3]{(x+3)}} }\) \(\LARGE\color{blue}{ [ 2(x+3)+1]\frac{1}{\sqrt[3]{(x+3)}} }\) \(\LARGE\color{blue}{ [ 2x+6+1]\frac{1}{\sqrt[3]{(x+3)}} }\) \(\LARGE\color{blue}{ (2x+7)\frac{1}{\sqrt[3]{(x+3)}} }\)

OpenStudy (anonymous):

Thanks @SolomonZelman